ICM 1998, Berlin, Aug. 18–27

Abstracts of Plenary and Invited Lectures

Section:

5. Differential Geometry and Global Analysis

1991 MS Classification: Joachim Lohkamp, Universität Augsburg, Germany Curvature Contents of Geometric Spaces

We discuss curvature relevant deformations of spaces. and indicate the existence of some individual capacity of a manifold (and more general spaces) measuring a maximal amount of curvature that could be carried by this space.

ICM 1998 Berlin · Plenary and Invited Lectures